OpenStudy (anonymous):

HEYY YALL! If your car masses 1000 kg, and you and your three friends (mass 60 kg each) get into the car, How much FORCE is on each tire? Can anyone show me how to do this question, and maybe include an answer so I could check my work? THANKS! :)

OpenStudy (amriju):

0

OpenStudy (anonymous):

why?

OpenStudy (amriju):

the car is not flying i believe....so no accelaration in vertical direction...rite...now all the forces u talked about ....the weights...they all act in vertical direction....but the car has no vertical accn...force must be 0.....accualy the normal force balances it all....

OpenStudy (amriju):

in which standard r u...natasha..?

OpenStudy (anonymous):

11th

OpenStudy (anonymous):

and thanks! :)

OpenStudy (amriju):

i get it .....accualy last day a kid of 7th standard asked a ques that required integration....god...

OpenStudy (anonymous):

oh wow!! can you help me with another ?

OpenStudy (anonymous):

4) If you weigh 800N and are on an elevator accelerating upwards at 2 m/s2, b. What Normal Force do you feel acting on you? c. Based on (b), how much do you feel that you weigh on that elevator?

OpenStudy (amriju):

look...here u r going up..rite..?...so ur net accelaration is 2 rite...but the weight is desparetely trying to make u go down rite...so the hero must be normal force..now we equate all forces N-800=ur mass*ur net accn... N the normal...must be more than the force due to gravity...and the net force acting on u must be ur mass *yr accn thats around 160newton...taking g=10... so we have N-800=160 thenN=960...answer may jst vary a little..coz i took g=10...and the weight u feel....

OpenStudy (amriju):

is of 960newton

OpenStudy (anonymous):

how did you get 160? shouldn't it be 1600?

OpenStudy (amriju):

sorry...i was jst out 4 awhile...

OpenStudy (amriju):

whats ur mass..? for the force on u ...u need ur mass* accelaration..ur mass is weight/g thats 800/10=80k's...DO NOT CONFUSE BETN MASS AND WEIGHT EVER...

OpenStudy (anonymous):

ohhh okay! i got it!

OpenStudy (amriju):

:D

OpenStudy (anonymous):

i have another question too! haah sorry but im going to make a new page so if you do answer it, i could give you best response! :)

OpenStudy (amriju):

kkkkkkkkkkk